Find the resultant and direction of 15N and 20N acting 75° with the horizontal



1
Expert's answer
2021-02-23T10:03:07-0500
"R=\sqrt{15^2+20^2+2(15)(20)\cos{75}}=27.9\ N\\\alpha=\arctan{\frac{20\sin{75}}{15+20\cos{75}}}=43.8\degree"